OpenStudy (anonymous):

evaluate f(x) when f(x)=7x-3.

myininaya (myininaya):

?

OpenStudy (anonymous):

do you mean when (x)=7x-3?

OpenStudy (anonymous):

maybe we have some numbers for x?

myininaya (myininaya):

i want to ride my bicycle! maybe solve f(x)=0 when f(x)=7x-3

OpenStudy (anonymous):

wats the exact question??????

myininaya (myininaya):

its a jeaporady question i believe. you have to guess

OpenStudy (anonymous):

f(x)=7x-3 when f(x)=7x-3 lol

myininaya (myininaya):

i guess she stepped out to smoke a cigarrate

OpenStudy (anonymous):

oh i am sorry typo error the problem is evaluate f(4) when f(x) =7x-3

OpenStudy (anonymous):

f(4) = 7(4) - 3 f(4) = 28 - 3 = 25

OpenStudy (anonymous):

U hv to replace x by 4 at all places and then calculate as i hv done